# Chemical reasoning in LLMs unlocks steerable synthesis planning and reaction mechanism elucidation

### v1

@@ 2.2 Mechanism Elucidation via LLM-Guided Search
A reaction mechanism is a specification of why and how a given chemical transformation occurs, by means of a set of elementary steps 43. The power of mechanisms in chemistry lies not only in its explanatory power of a single reaction instance, but also in that the reach of an explanation may extend further than only that reaction; potentially explaining more observed reactions, and also even predicting potential unknown transformations 42.
Figure 4: a) Requirements and impact of mechanistic elucidation. b) Example of actions in our mechanism framework displayed on example structures, with a post-processing interpretation. c) Example task broken down into moves compliant with our mechanism game framework. d) Performance of the 5 models tested averaged over each task. e) Comparison of global performance with and without expert prompt guidance. f) All starting reactants and goal products of our 12 tasks, grouped by category.
We define a set of elementary steps that can be applied to any given molecular set (Methods 4.4 and Figure 4b); these elementary steps can generally be applied to any bond/atom in any molecule, and thus serve as a fundamental basis for formulating mechanisms. The task is then to, given a chemical reaction, find a suitable sequence of such elementary steps that connects the reactants with the products. The search for the best path between reactants and products in chemistry has already been tackled using different ML paradigms, whether using generative models in rule-based environments 8, contrastive learning 13, reinforcement learning in 3D space 53 or flow-matching 14. In this work, we show that LLMs reasoning can be useful for this, even in complex cases. In our approach, an LLM analyzes a partially constructed solution and rates a proposed mechanistic step that continues such solution. For evaluation, we designed a benchmark comprising 12 diverse reactions (Figure 4f) along with their mechanisms (SI-D). Performance is measured as the relative score gap between ground truth and alternative moves, averaged over all the steps in the mechanism (Methods 4.6). A perfect scoring model would highly rate correct over incorrect moves, thus this metric correlates with performance in a real search setting by assessing the selectivity of the model at each step.
@@
Task # moves GPT-4o Mini GPT-4o DeepSeek V3 Claude 3.5 Claude 3.7
@@
1 4 $50\pm 16$ $55\pm 19$ $75\pm 0$ $90\pm 12$ $\mathbf{100\pm 0}$
2 4 $70\pm 10$ $60\pm 12$ $70\pm 10$ $95\pm 10$ $\mathbf{100\pm 0}$
3 4 $55\pm 10$ $70\pm 10$ $80\pm 10$ $75\pm 0$ $\mathbf{100\pm 0}$
4 4 $30\pm 10$ $50\pm 16$ $55\pm 10$ $\mathbf{75\pm 0}$ $\mathbf{75\pm 0}$
5 4 $35\pm 12$ $50\pm 0$ $50\pm 16$ $\mathbf{80\pm 10}$ $\mathbf{80\pm 10}$
6 8 $25\pm 8$ $52\pm 5$ $70\pm 13$ $80\pm 6$ $\mathbf{85\pm 5}$
7 12 $20\pm 10$ $48\pm 9$ $20\pm 15$ $75\pm 0$ $\mathbf{80\pm 6}$
8 8 $5\pm 4$ $18\pm 10$ $28\pm 7$ $\mathbf{68\pm 3}$ $\mathbf{68\pm 10}$
9 4 $30\pm 10$ $30\pm 10$ $60\pm 25$ $30\pm 10$ $\mathbf{80\pm 10}$
10 14 $24\pm 10$ $43\pm 6$ $36\pm 12$ $66\pm 3$ $\mathbf{79\pm 5}$
11 8 $32\pm 6$ $28\pm 9$ $35\pm 5$ $50\pm 8$ $\mathbf{62\pm 8}$
12 8 $22\pm 9$ $18\pm 6$ $18\pm 6$ $18\pm 6$ $\mathbf{42\pm 6}$
Table 1: Percentage of ground truth moves classified as strict top-1 when scored with 5 alternatives at each step. These scores are reported for LLMs alone, without the help of an expert prompt defining the mechanism.
A key advantage of search guided by an LLM is that it allows the specification of any arbitrary amount of information, parameters and instructions, through the LLM’s text interface. Particularly relevant for this task, the input consists of the reaction (reactants and products), but could also contain the reaction conditions (solvent, temperature, concentrations, etc), any available experimental data (such as kinetic studies), among others. In this line of ideas, we experiment with text-guided search, where search is guided by an external text describing the sequence of steps in a hypothetical mechanism. The search here thus functions as a decoder from the input text+reaction, into a proper sequence of elementary steps. The source of this external text can be an expert human, or another LLM which, as is clear from the previous sections, can accurately analyze reactions and describe their mechanisms. Figure 4e shows that adding such information is generally beneficial for performance, boosting the performances of poorly performing LLMs, as is the case of gpt-4o-mini and gpt-4o. Notably, Claude-3.7’s base performance is already higher than any other competitor, with expert prompts bringing additional but less substantial improvements. These results show the potential for including multiple sources of information as input, hinting at the future possibility of automated mechanistic (hypothesis) generation and closed-loop refinement.
@@ 3 Discussion
In this work, we demonstrate that LLMs can serve as powerful chemical reasoning engines and present a framework to exploit these capabilities in scientifically relevant tasks in synthetic chemistry. Our approach combines traditional search algorithms with LLMs’ sophisticated reasoning abilities, positioning the models as expert evaluators that guide search toward solutions that are aligned with chemical principles as well as expert intent. This integration helps overcome fundamental limitations in existing tools while enabling more intuitive interfaces for complex chemical tasks.
@@
Our results first establish that current LLMs possess remarkable capabilities for detailed chemical reasoning, accurate analysis and evaluation of chemical objects at multiple scales, from functional groups in molecules to strategic patterns in full synthetic routes. We leverage these capabilities through a framework where potential solutions are generated by traditionally used computational environments, while LLMs analyze intermediate solutions to assess their validity towards constructing the desired solution.
We showcase the practical utility of this approach through two challenging applications: steerable synthesis planning and mechanism elucidation. In synthesis planning, our system enables natural language specification of strategic constraints, allowing chemists to guide Computer Assisted Synthesis Planning (CASP) searches towards solutions that are more aligned with their intent. We furthermore evaluate the system on experimentally validated routes for complex targets like Atorvastatin, and show how the system correctly identifies and evaluates key strategic elements, not only providing key insights into the analyzed solutions but also validating the practical potential of this approach.
@@
Similarly for mechanism elucidation, we show that a similar approach can successfully guide search towards chemically feasible mechanisms, starting only from a set of simple and general elementary steps, making our approach generally applicable to a large space of organic reactions. We note that any arbitrary specification, guidance, or initial condition can be encoded as input to the system, thanks to the flexibility of the LLMs’ text interface. We show the potential of this idea by experimenting with "expert guidance" prompts in the input, which resemble intuitions of expert chemists about how the mechanism of a given reaction should look like. These considerations can be extended to include feedback from experiments, that iteratively refine mechanistic hypotheses.
However key limitations remain. On our most complex synthetic tasks (Strychnine with 26 steps), LLMs exhibited a number of failure modes, some of which can be traced back to the several stages of pre- and post-training, that bias models’ responses towards shorter and overly simplistic and optimistic responses. This largely limits the analysis and application of our framework to larger routes and thus more complex targets. Some of these limitations can be tackled through improved and more specific fine-tuning techniques, data collection methods, among other techniques that better align the models with the tasks presented here. Still, as we show in this work, commercially available models are already very strong baselines on these tasks, which highlights the potential for future improvements.
@@ 4.1 Large Language Models
Throughout this work we used several LLMs from different providers, all through litellm 54 as it provides an unified interface to multiple provider’s APIs. A temperature of 0.1 has been used across all LLMs. The specific model versions used throughout this work are as follows: gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, claude-3-5-sonnet-20241022, claude-3-7-sonnet-20250219.
4.2 Steerable synthesis planning
@@
In this work we present a framework for steerable synthesis planning, which integrates LLMs as evaluators within a traditional retrosynthetic search algorithm. The process begins with a target molecule and a natural language query specifying desired strategic features of the synthetic route. We employed the AiZynthfinder software 16 as the underlying retrosynthesis engine to generate a diverse set of potential synthetic routes for a given target molecule. For each candidate route generated by the underlying retrosynthesis software, we constructed a detailed textual representation of the route, including SMILES strings of all intermediates and reactants, and the sequence of transformations applied at each step. This textual representation, along with the user-provided natural language query, is then passed to the LLM. The LLM is prompted to analyze the synthetic route and evaluate its alignment with the strategic requirements specified in the query. The LLM’s output is a score reflecting the degree to which the route satisfies the user’s query, along with a textual rationale justifying its assessment. This score is then used to rank and filter the candidate routes generated by AiZynthfinder, prioritizing routes that are deemed strategically relevant by the LLM. This process allows us to leverage the systematic exploration capabilities of AiZynth while incorporating the sophisticated chemical reasoning of LLMs to guide the search towards strategically desirable synthetic pathways.
4.3 Steerable synthesis planning - Benchmark
@@
To evaluate the performance of our steerable synthesis planning framework, we created a benchmark consisting of pairs of molecular targets and steering prompts. We selected a diverse set of molecular targets that represent various levels of complexity and synthetic challenges. These targets include both well-studied molecules and novel compounds to ensure a comprehensive evaluation. For each molecular target, we crafted steering prompts that specified desired strategic elements in the synthetic route. These prompts ranged from simple reaction preferences to complex strategic requirements. The specific prompts used are given in the SI B.
@@
As the system returns a set of routes, together with alignment scores, the primary evaluation metric reported in this work is correlation with ground truth scores. The scores are computed using scripts tailored for each target-prompt pair, and work by finding a specifically defined event in a synthetic route. For example, if the prompt requires "an early ring formation", then this is translated into a script that finds ring-forming reactions, together with the relative position of this reaction, if any, in the route. A score is thus calculated to account for the happening of the reaction, and it’s relative position, in this case giving a higher score to routes where a ring-forming reaction occurs in the early stages (far from the final product), and progressively lower scores for routes when this happens in later stages. The minimum score is given to routes where the ring-forming reaction doesn’t happen. The full scripts can be found in the github repository.
@@ 4.5 Mechanism elucidation - Benchmark
The mechanism elucidation benchmark was designed to include diverse cases of chemical reactions. It includes several reaction types as well as different scales of molecular weight of the products, assessing different levels of understanding of chemical principles, molecular and reaction representations, and ability to understand the rules proposed by the environment. In Figure 4c and SI-D, it is shown how a mechanism is broken down into ionizations and attacks only.

### v2

@@ 2.3 Mechanism Elucidation via LLM-Guided Search
A reaction mechanism is a specification of why and how a given chemical transformation occurs, by means of a set of elementary steps [43]. The power of mechanisms in chemistry lies not only in its explanatory power of a single reaction instance, but also in that the reach of an explanation may extend further than only that reaction; potentially explaining more observed reactions, and also even predicting potential unknown transformations [42].
Figure 4: Mechanisms elucidation a) Requirements and impact of mechanistic elucidation. b) Example of actions in our mechanism framework displayed on example structures, with a post-processing interpretation. c) Example task broken down into moves compliant with our mechanism game framework. d) Model separation performance, averaged across the full mechanism in each case, with n=5 repetitions. e) Comparison of global performance with and without guidance prompts. f) All starting reactants and goal products of our 12 tasks, grouped by category.
We define a set of elementary steps that can be applied to any given molecular set (Methods 4.5 and Figure 4b); these elementary steps can generally be applied to any bond/atom in any molecule, and thus serve as a fundamental basis for formulating mechanisms. The task is then to, given a chemical reaction, find a suitable sequence of such elementary steps that connects the reactants with the products. The search for the best path between reactants and products in chemistry has already been tackled using different ML paradigms, whether using generative models in rule-based environments [8], contrastive learning [13], reinforcement learning in 3D space [57] or flow-matching [14]. Here, we show that LLMs reasoning can be useful for this, even in complex cases. In our approach, an LLM analyzes a partially constructed solution and rates a proposed mechanistic step that continues such solution. For evaluation, we designed a benchmark comprising 12 diverse reactions (Figure 4f) along with their mechanisms (SI-LABEL:si:mechbench-rxns). Performance is measured as the relative score gap between ground truth and alternative moves, averaged over all the steps in the mechanism (Methods 4.7). A perfect scoring model would highly rate correct over incorrect moves, thus this metric correlates with performance in a real search setting by assessing the selectivity of the model at each step.
@@
Task (# steps) 4o Mini 4o DS-3 Cd 3.5 Cd 3.7 Gm 2.5
@@
1 (4) $50\pm 16$ $55\pm 19$ $75\pm 0$ $90\pm 12$ $\mathbf{100\pm 0}$ $\mathbf{100\pm 0}$
2 (4) $70\pm 10$ $60\pm 12$ $70\pm 10$ $95\pm 10$ $\mathbf{100\pm 0}$ $95\pm 10$
3 (4) $55\pm 10$ $70\pm 10$ $80\pm 10$ $75\pm 0$ $\mathbf{100\pm 0}$ $\mathbf{100\pm 0}$
4 (4) $30\pm 10$ $50\pm 16$ $55\pm 10$ $75\pm 0$ $75\pm 0$ $\mathbf{90\pm 12}$
5 (4) $35\pm 12$ $50\pm 0$ $50\pm 16$ $\mathbf{80\pm 10}$ $\mathbf{80\pm 10}$ $70\pm 19$
6 (8) $25\pm 8$ $52\pm 5$ $70\pm 13$ $80\pm 6$ $\mathbf{85\pm 5}$ $78\pm 39$
7 (12) $20\pm 10$ $48\pm 9$ $20\pm 15$ $75\pm 0$ $\mathbf{80\pm 6}$ $72\pm 37$
8 (8) $5\pm 4$ $18\pm 10$ $28\pm 7$ $\mathbf{68\pm 3}$ $\mathbf{68\pm 10}$ $67\pm 34$
9 (4) $30\pm 10$ $30\pm 10$ $60\pm 25$ $30\pm 10$ $\mathbf{80\pm 10}$ $40\pm 30$
10 (14) $24\pm 10$ $43\pm 6$ $36\pm 12$ $66\pm 3$ $\mathbf{79\pm 5}$ $64\pm 23$
11 (8) $32\pm 6$ $28\pm 9$ $35\pm 5$ $50\pm 8$ $\mathbf{62\pm 8}$ $57\pm 30$
12 (8) $22\pm 9$ $18\pm 6$ $18\pm 6$ $18\pm 6$ $42\pm 6$ $\mathbf{55\pm 19}$
Table 1: Top-1 accuracy on mechanisms elucidation tasks Percentage of ground truth moves classified as strict top-1 when scored with 5 alternatives at each step. Each row displays the results for each of the tasks on the left column. The number of moves involved in each task are specified in parentheses. The reported scores are percentage of correctly scored moves. Scores are reported for LLMs alone, without the help of a guidance prompt defining the mechanism. Each column displays the results for each LLM. The column names used stand for: 4o-mini: GPT-4o-mini, 4o: GPT-4o, DS-3: DeepSeek-3, Cd 3.5: Claude-3.5-Sonnet, Cd 3.7: Claude-3.7-Sonnet, Gm 2.5: Gemini-2.5-pro.
A key advantage of search guided by an LLM is that it allows the specification of any arbitrary amount of information, parameters and instructions, through the LLM’s text interface. Particularly relevant for this task, the input consists of the reaction (reactants and products), but could also contain the reaction conditions (solvent, temperature, concentrations, etc), any available experimental data (such as kinetic studies), among others. In this line of ideas, we experiment with text-guided search, where search is guided by an external text describing the sequence of steps in a hypothetical mechanism. The search here thus functions as a decoder from the input text+reaction, into a proper sequence of elementary steps. The source of this external text can be an expert human, or another LLM which, as is clear from the previous sections, can accurately analyze reactions and describe their mechanisms. Figure 4e shows that adding such information is generally beneficial for performance, boosting the performances of poorly performing LLMs, as is the case of gpt-4o-mini and gpt-4o. Notably, even though Gemini-2.5-Pro’s base performance is already higher than any other competitor, guidance prompts still provide additional and substantial improvements. These results show the potential for including multiple sources of information as input, hinting at the future possibility of automated mechanistic (hypothesis) generation and closed-loop refinement.
@@ 3 Discussion
In this work, we demonstrate that LLMs can serve as powerful tools enabling chemical analysis at arbitrary levels of abstraction. We present Synthegy: a framework to exploit these capabilities in scientifically relevant tasks in synthetic chemistry. Our approach combines traditional search algorithms with LLMs’ sophisticated reasoning abilities, positioning the models as expert evaluators that can filter, rerank, or guide search toward solutions that are aligned with chemical principles as well as expert intent, across a wide rang of chemical scales. This integration helps overcome fundamental limitations in existing tools while enabling more intuitive interfaces for complex chemical tasks.
@@
Our results first establish that current LLMs possess remarkable capabilities for detailed chemical reasoning, accurate analysis and evaluation of chemical objects at multiple scales, from functional groups in molecules, to single reactions, to strategic patterns in full synthetic routes. We leverage these capabilities through a framework where candidate solutions are generated by traditionally used computational environments, while LLMs analyze proposed solutions to assess their validity in relation to the specified target solution. Furthermore, these capabilities can be leveraged to directly guide search algorithms towards desirable and chemically-sensible solutions.
We showcase the practical utility of this approach through two challenging applications: strategy-aware synthesis planning and mechanism elucidation. In synthesis planning, our system enables natural language specification of strategic constraints, allowing chemists to select solutions from computer assisted synthesis planning algorithms, towards routes that are more aligned with their intent. Such intent can range from specifications of the reaction sequences used or relative timings of scaffold formation, to full assessments of general feasibility, enabling a novel way of selecting promising and experimentally feasible synthetic routes. We demonstrate this through an experimental study where Synthegy successfully selects the experimentally validated route as highly-feasible, from a non-exhaustive pool of candidate solutions proposed by several standard retrosynthesis tools.
We furthermore evaluate the system on experimentally validated routes for complex targets like Strychnine, and show how the system correctly identifies and evaluates key strategic elements, not only providing key insights into the analyzed solutions but also further validating the practical potential of Synthegy to select solutions with specific properties, even for complex molecular scaffolds and strategic requirements.
@@
Similarly for mechanism elucidation, we show that a similar approach can successfully guide search towards chemically feasible mechanisms, starting only from a set of simple and general elementary steps, making our approach generally applicable to a large space of organic reactions. We note that any arbitrary specification, guidance, or initial condition can be encoded as input to the system, thanks to the flexibility of the LLMs’ text interface. We show the potential of this idea by experimenting with ”guidance prompts” in the input, which resemble intuitions of expert chemists about how the mechanism of a given reaction should look like. These considerations can be extended to include feedback from experiments, that iteratively refine mechanistic hypotheses.
However, some limitations remain. LLMs exhibited a number of failure modes, more prominently eventual misunderstandings of the format of the input data, which drives them to analyze reactions in an inverse sense and thus leading them to deem routes unfeasible. Furthermore, some models display lazyness, which bias models’ responses towards shorter and overly simplistic and optimistic responses. Such behavior can be traced back to pre- and post-training techniques used to train these models. These limitations can be tackled by improved prompting techniques, task-specific fine-tuning, among others. Even with these limitations, we show that the current models are very strong baselines, highlighting the potential for future improvements. Moreover, we show that success in the current implementation is very much dependent on the quality and diversity of synthetic routes generated by the backend retrosynthesis systems. Further research needs to go into the application of these strong capabilities for directly guiding search methods towards solutions that align with prompts and chemical principles, overcoming the inherent limitations of predefined sets of solutions.
@@ 4.1 Large Language Models
Throughout this work we used several LLMs from different providers, all through litellm [58] as it provides a unified interface to multiple provider’s APIs. A temperature of 0.1 has been used across all LLMs. The specific models used throughout this work include: OpenAI models [49] (gpt-4o, gpt-4o-mini, gpt-4-turbo, o3, o4-mini), Anthropic Claude models [59] (claude-3-sonnet-20240229, claude-3-opus-20240229, claude-3-5-sonnet-20241022, claude-3-7-sonnet-20250219, claude-opus-4-20250514, claude-4-sonnet-20250522, claude-2.1) , Meta LLaMA models [60] (llama-3.3-70b-instruct, llama-3.3-8b-instruct, llama-4-scout, llama-4-maverick, llama-3.1-405b-instruct), DeepSeek models [50] (deepseek-r1-0528, deepseek-chat-v3-0324, deepseek-r1-distill variants), Google Gemini models [51] (gemini-2.5-pro-preview, gemini-flash-1.5, gemini-pro-1.5, gemini-2.0-flash-001, gemini-2.5-flash-preview), and Qwen [61] (qwen3-32b).
4.2 Strategy-aware synthesis planning
@@
In this work we present a framework for strategy-aware synthesis planning, which integrates LLMs as evaluators within a traditional retrosynthetic search algorithm. The process begins with a target molecule and a natural language query specifying desired strategic features of the synthetic route. We employed the AiZynthfinder software [16] as the underlying retrosynthesis engine to generate a diverse set of potential synthetic routes for a given target molecule. For each candidate route generated by the underlying retrosynthesis software, we constructed a detailed textual representation of the route, including SMILES strings of all intermediates and reactants, and the sequence of transformations applied at each step. This textual representation, along with the user-provided natural language query, is then passed to the LLM. The LLM is prompted to analyze the synthetic route and evaluate its alignment with the strategic requirements specified in the query. The LLM’s output is a score reflecting the degree to which the route satisfies the user’s query, along with a textual rationale justifying its assessment. This score is then used to rank and filter the candidate routes generated by AiZynthfinder, prioritizing routes that are deemed strategically relevant by the LLM. This process allows us to leverage the systematic exploration capabilities of AiZynth while incorporating the sophisticated chemical reasoning of LLMs to guide the search towards strategically desirable synthetic pathways.
4.3 Feasibility assessment
To evaluate the capability of Synthegy to assess synthetic route feasibility, we collected synthetic routes from multiple sources and compared their LLM-assigned feasibility scores. Routes were obtained from three different retrosynthesis engines alongside experimentally validated routes from the literature.
We accessed synthetic routes through the following platforms:
AiZynth: Synthetic routes were obtained using the AiZynth retrosynthesis planning software [16] with default configuration parameters. The software was run locally using standard installation procedures and default model weights.
Reaxys: Synthetic routes were obtained through the Reaxys API using default parameters as specified by the provider. Access was granted through institutional licensing agreements. The API was queried using target molecule SMILES strings, and the top-ranked routes (by Reaxys scoring) were collected for each target.
Synthia: Routes were generated using Synthia-Lite through their web interface with default parameter settings. No specific parameter customization was performed, utilizing the platform’s standard search algorithms and scoring functions. Routes were collected in SMILES format for subsequent analysis.
Experimental Routes: Literature routes for targets T1-T4 were manually extracted from their respective publications [54, 55, 56] and converted to SMILES format following the same reaction representation standards used for computationally generated routes.
Feasibility Scoring:
All collected routes were evaluated using the Synthegy framework with gemini-2.5-pro as the backend LLM, selected based on its superior performance in our benchmark evaluation (Section 2.1). The feasibility assessment prompt (Section 2.2) was applied consistently across all route sources. Each route was scored independently, with the LLM providing both numerical scores (0-10 scale) and detailed rationales for feasibility assessments.
4.4 Strategy-aware synthesis planning - Benchmark
@@
To evaluate the performance of our strategy-aware synthesis planning framework, we created a benchmark consisting of pairs of molecular targets and prompts. We selected a diverse set of molecular targets that represent various levels of complexity and synthetic challenges. These targets include both well-studied molecules and novel compounds to ensure a comprehensive evaluation. For each molecular target, we crafted prompts that specified desired strategic elements in the synthetic route. These prompts ranged from simple reaction preferences to complex strategic requirements. The specific prompts used are given in the SI LABEL:si:steering-prompts.
@@
As the system returns a set of routes, together with alignment scores, the primary evaluation metric reported in this work is correlation with ground truth scores. The scores are computed using scripts tailored for each target-prompt pair, and work by finding a specifically defined event in a synthetic route. For example, if the prompt requires ”an early ring formation”, then this is translated into a script that finds ring-forming reactions, together with the relative position of this reaction, if any, in the route. A score is thus calculated to account for the happening of the reaction, and it’s relative position, in this case giving a higher score to routes where a ring-forming reaction occurs in the early stages (far from the final product), and progressively lower scores for routes when this happens in later stages. The minimum score is given to routes where the ring-forming reaction doesn’t happen. The full scripts can be found in the github repository.
@@ 4.6 Mechanism elucidation - Benchmark
The mechanism elucidation benchmark was designed to include diverse cases of chemical reactions. It includes several reaction types as well as different scales of molecular weight of the products, assessing different levels of understanding of chemical principles, molecular and reaction representations, and ability to understand the rules proposed by the environment. In Figure 4c and SI-LABEL:si:mechbench-rxns, it is shown how a mechanism is broken down into ionizations and attacks only.